# Extract and Diffuse: Latent Integration for Improved Diffusion-based Speech and Vocal Enhancement

### v1

@@ Abstract
Diffusion-based generative models have recently achieved remarkable results in speech and vocal enhancement due to their ability to model complex speech data distributions. While these models generalize well to unseen acoustic environments, they may not achieve the same level of fidelity as the discriminative models specifically trained to enhance particular acoustic conditions. In this paper, we propose Ex-Diff, a novel score-based diffusion model that integrates the latent representations produced by a discriminative model to improve speech and vocal enhancement, which combines the strengths of both generative and discriminative models. Experimental results on the widely used MUSDB dataset show relative improvements of 3.7% in SI-SDR and 10.0% in SI-SIR compared to the baseline diffusion model for speech and vocal enhancement tasks, respectively. Additionally, case studies are provided to further illustrate and analyze the complementary nature of generative and discriminative models in this context.
I Introduction
@@
In this paper, we propose Extract and Diffuse (Ex-Diff), a novel approach that combines the strengths of both generative and discriminative methods for improved diffusion-based speech and vocal enhancement. Ex-Diff leverages pre-trained latent representations from the universal source separation (USS) model [9], a discriminative model combining sound event detection (SED) with a conditional source separation model, as conditioning inputs to a score-based diffusion model via a cross-attention mechanism. Unlike the approach in [28], which conditions directly on the audio mixture, using latent representations provides a clearer indication of what to extract and enhance, thus reducing the likelihood of generating unexpected sounds while benefiting from the improved quality of generative modeling. We evaluate Ex-Diff on the VoiceBank-DEMAND [30] and MUSDB18 [31] datasets. The results show Ex-Diff achieves similar SI-SDR on VoiceBank-DEMAND and a 3.7% relative improvement in SI-SDR on MUSDB18 compared to the baseline diffusion model [28]. [^1] The demo can be accessed at: https://liuzhan22.github.io/ex-diff-demo/.
II Related Work
II-A Discriminative Models for Speech and Vocal Enhancement
@@
II-B Diffusion Models for Speech and Vocal Enhancement
@@
III Diffusion for Speech Enhancement
@@ III-2 Reverse Process
$\mathrm{d}\mathbf{x}_{t}=\left[g\left(t\right)^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}\left(\mathbf{x}_{t}\mid\mathbf{y}\right)-f\left(\mathbf{x}_{t},\mathbf{y}\right)\right]\mathrm{d}t$ (3)
$+g\left(t\right)\mathrm{d}\bar{\mathbf{w}}$
@@
III-3 Training Objective
This section discusses the objective function used to train $s_{\theta}(\mathbf{x}_{t},\mathbf{y},t)$. Since Eqn. (1) defines a Gaussian process, the mean and variance of step $\mathbf{x}_{t}$ are easily determined given the initial state $\mathbf{x}_{0}$ and condition $\mathbf{y}$ [35]:
@@ III-3 Training Objective
$\sigma(t)^{2}=\frac{\sigma_{\mathrm{min}}^{2}\left(\left(\sigma_{\mathrm{max}}/\sigma_{\mathrm{min}}\right)^{2t}-\mathrm{e}^{-2\gamma t}\right)\log\left(\sigma_{\mathrm{max}}/\sigma_{\mathrm{min}}\right)}{\gamma+\log\left(\sigma_{\mathrm{max}}/\sigma_{\mathrm{min}}\right)}.$
@@ IV Extract and Diffuse
IV-A Latent Representation Extraction
Following the procedure in USS [9], a pre-trained audio neural networks (PANNs) model [37] is used to extract the latent representation of vocals in the input audio. The PANNs model is trained with weakly labeled data from AudioSet [38], and it can localize the occurrence of sound classes as a SED system. For the audio clip, the SED system can produce a frame-wise event prediction: $p_{\text{SED}}\in[0,1]^{T\times K}$, where $T$ is the number of frames and $K$ is the predefined number of sound classes. In our model, the output of the penultimate layer of this PANNs model is used with a size of $(T,H)$, and its averaged vector with a size of $H$ is obtained as the latent representation $\mathbf{l}=\frac{1}{T}\sum\nolimits_{t=1}^{T}\mathbf{h}_{t}$, where $\mathbf{h}_{t}\in\mathbb{R}^{H}$ is the feature vector of the $t$-th frame. The default value of $H$ is set to 2048.
@@ IV-A Latent Representation Extraction
IV-B The Conditional Reverse Diffusion Process in Ex-Diff
@@
V Experimental Setup
V-A Datasets and Evaluation Metrics
@@
TABLE I: Left: Speech enhancement results obtained for VoiceBank-DEMAND test set. Right: Vocal enhancement results obtained for the mix3dB-MUSDB test set. Each result is presented in the form of mean $\pm$ standard deviation. AS, M3 and VB refer to the AudioSet, MUSDB-mix3dB and VB-DMD datasets, respectively.
@@
USS [9] AS $2.19\pm 0.56$ $0.79\pm 0.14$ $16.8\pm 4.8$ $31.0\pm 11.6$ $17.8\pm 4.7$ $1.25\pm 0.31$ $0.29\pm 0.20$ $-2.5\pm 6.0$ $28.8\pm 8.9$ $-2.5\pm 6.0$
USS+SGMSE AS+VB / M3 $2.59\pm 0.61$ $0.83\pm 0.12$ $\textbf{17.9}\pm\textbf{4.7}$ $\textbf{37.4}\pm\textbf{9.9}$ $\textbf{18.1}\pm\textbf{4.7}$ $1.24\pm 0.35$ $0.25\pm 0.20$ $-3.8\pm 5.7$ $\textbf{29.5}\pm\textbf{8.2}$ $-3.8\pm 5.7$
@@ V-B Training Details
VI Experimental Results
VI-A Speech Enhancement Results
The left part of Table I presents the speech enhancement results on the VoiceBank-DEMAND test set. Our proposed Ex-Diff model shows strong performance across a range of objective metrics, showcasing significant improvements over USS in PESQ, ESTOI, and SI-SDR, achieving similar performance in SGMSE, and surpassing it in the SI-SIR metric. However, it is worth noting that the combined USS+SGMSE model exhibits a marginally better SI-SDR score compared to Ex-Diff, which can be attributed to two key factors. First, USS as a discriminative model, uses an L1 loss function directly calculated on the predicted and ground truth waveforms, inherently favouring higher SI-SDR scores due to the better (compared to the diffusion training loss) correlation between minimizing waveform distance and maximizing signal-to-interference ratio [45, 46]. Second, the subsequent refinement using the SGMSE diffusion model further contributes to this bias by directly targeting waveform similarity, a factor heavily reflected in SI-SDR. Conversely, Ex-Diff, as an end-to-end generative model, focuses on learning the underlying speech distribution to produce enhanced speech with high perceptual quality. While this approach may not prioritize SI-SDR optimization to the same extent as a concatenated discriminative-diffusion approach, it avoids potential drawbacks of model cascading, such as error propagation and increased complexity. Furthermore, it is crucial to recognize that SI-SDR alone cannot fully encapsulate the nuanced aspects of speech quality, such as naturalness and listener preference, areas where generative approaches like Ex-Diff often excel.
VI-B Vocal Enhancement Results
The MUSDB18 dataset is used for vocal enhancement. According to [47], training models with an appropriate SNR and using incoherent mixing for data augmentation can lead to better performance. We reprocessed the SNR of the MUSDB18 dataset to 3dB, and adopted an incoherent mixing strategy for the training set, which enhances both the quantity and quality of the data. The experimental results on MUSDB-mix3dB test set, as shown in Table I, indicate that our model outperforms all baseline models. It should be noted that, to ensure a fair comparison, our baseline models were also trained from scratch on the MUSDB-mix3dB dataset. In which the USS model achieves good results on the original MUSDB dataset, but performs poorly on the processed MUSDB dataset, which also reflects that the generative model has stronger generalization ability.
VI-C Discussions on Selected Examples
@@ VI-D Ablation Study
TABLE II: Results of the ablation study. A single cross-attention layer (as in Fig. 1) performs the best. Removing the attention mechanism by concatenating the latent representation with $\mathbf{x}$ and $\mathbf{y}$, using a full Transformer-like block, or adding up to three attention blocks did not yield better results.

### v2

@@ Abstract
Diffusion-based generative models have recently achieved remarkable results in speech and vocal enhancement due to their ability to model complex speech data distributions. While these models generalize well to unseen acoustic environments, they may not achieve the same level of fidelity as the discriminative models specifically trained to enhance particular acoustic conditions. In this paper, we propose Ex-Diff, a novel score-based diffusion model that integrates the latent representations produced by a discriminative model to improve speech and vocal enhancement, which combines the strengths of both generative and discriminative models. Experimental results on the widely used MUSDB dataset show relative improvements of 3.7% in SI-SDR and 10.0% in SI-SIR for vocal enhancement tasks compared to the baseline diffusion model. Additionally, case studies are provided to further illustrate and analyze the complementary nature of generative and discriminative models in this context.
1 Introduction
@@
In this paper, we propose Extract and Diffuse (Ex-Diff), a novel approach that combines the strengths of both generative and discriminative methods for improved diffusion-based speech and vocal enhancement. Ex-Diff leverages pre-trained latent representations from the universal source separation (USS) model [23], a discriminative model combining sound event detection (SED) with a conditional source separation model, as conditioning inputs to a score-based diffusion model via a cross-attention mechanism. Unlike the approach in [32], which conditions directly on the audio mixture, using latent representations provides a clearer indication of what to extract and enhance, thus reducing the likelihood of generating unexpected sounds while benefiting from the improved quality of generative modeling. We evaluate Ex-Diff on the VoiceBank-DEMAND [40] and MUSDB18 [30] datasets. The results show Ex-Diff achieves similar SI-SDR on VoiceBank-DEMAND and a 3.7% relative improvement in SI-SDR on MUSDB18 compared to the baseline diffusion model [32].
2 Related Work
2.1 Discriminative Models for Speech and Vocal Enhancement
@@
2.2 Diffusion Models for Speech and Vocal Enhancement
@@
3 Diffusion for Speech Enhancement
@@ 3.0.2 Reverse Process
$\mathrm{d}\mathbf{x}_{t}=\left[g\left(t\right)^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}\left(\mathbf{x}_{t}\mid\mathbf{y}\right)-f\left(\mathbf{x}_{t},\mathbf{y}\right)\right]\mathrm{d}t+g\left(t\right)\mathrm{d}\bar{\mathbf{w}}$ (3)
@@
3.0.3 Training Objective
This section discusses the objective function used to train $s_{\theta}(\mathbf{x}_{t},\mathbf{y},t)$. This section discusses the objective function used to train $s_{\theta}(\mathbf{x}_{t},\mathbf{y},t)$. Since Eqn. (1) defines a Gaussian process, the mean and variance of step $\mathbf{x}_{t}$ are easily determined given the initial state $\mathbf{x}_{0}$ and condition $\mathbf{y}$ [39]:
@@ 3.0.3 Training Objective
$\sigma(t)^{2}=\frac{\sigma_{\mathrm{min}}^{2}\left(\left(\sigma_{\mathrm{max}}/\sigma_{\mathrm{min}}\right)^{2t}-\mathrm{e}^{-2\gamma t}\right)\log\left(\sigma_{\mathrm{max}}/\sigma_{\mathrm{min}}\right)}{\gamma+\log\left(\sigma_{\mathrm{max}}/\sigma_{\mathrm{min}}\right)}.\$
@@ 4 Extract and Diffuse
4.1 Latent Representation Extraction
Following the procedure in USS [23], a pre-trained audio neural networks (PANNs) model [21] is used to extract the latent representation of vocals in the input audio.
The PANNs model is trained with weakly labeled data from AudioSet [9], and it can localize the occurrence of sound classes as a SED system. For the audio clip, the SED system can produce a frame-wise event prediction $p_{\text{SED}}\in[0,1]^{T\times K}$, where $T$ is the number of frames and $K$ is the predefined number of sound classes. In our model, each raw audio clip is first divided into segments of 64,000 samples. For the $i$-th segment, the PANNs model yields penultimate-layer features of size $(T,H)$, which are averaged over the temporal dimension to produce a segment-level vector $\mathbf{l}_{i}=\tfrac{1}{T}\sum\nolimits_{t=1}^{T}\mathbf{h}_{t}^{(i)}$, where $\mathbf{h}_{t}^{(i)}\in\mathbb{R}^{H}$ denotes the feature of the $t$-th frame in segment $i$. The segment-level vectors are then concatenated in temporal order to form the final latent representation $\mathbf{l}$ with shape $(L,H)$, where $L$ is the number of segments and the default dimension is H=2048H{=}2048. To preserve temporal dependencies, positional encodings are added, and cross-attention is applied along the temporal dimension.
@@ 4.1 Latent Representation Extraction
4.2 The Conditional Reverse Diffusion Process in Ex-Diff
@@
5 Experimental Setup
5.1 Datasets and Evaluation Metrics
@@
Table 1: Left: Speech enhancement results obtained for VoiceBank-DEMAND test set. Right: Vocal enhancement results obtained for the mix3dB-MUSDB test set. Note that models trained on VB or AS+VB are evaluated only on the speech enhancement task, while models trained on M3 are evaluated only on the vocal enhancement task. Each result is presented in the form of mean $\pm$ standard deviation. AS, M3 and VB refer to the AudioSet, MUSDB-mix3dB and VB-DMD datasets, respectively.
@@
USS [23] AS $2.19\pm 0.56$ $0.79\pm 0.14$ $16.8\pm 4.8$ $31.0\pm 11.6$ $17.8\pm 4.7$ $1.25\pm 0.31$ $0.29\pm 0.20$ $2.5\pm 6.0$ $28.8\pm 8.9$ $2.5\pm 6.0$
USS+SGMSE AS+VB / M3 $2.59\pm 0.61$ $0.83\pm 0.12$ $\textbf{17.9}\pm\textbf{4.7}$ $\textbf{37.4}\pm\textbf{9.9}$ $\textbf{18.1}\pm\textbf{4.7}$ $1.24\pm 0.35$ $0.25\pm 0.20$ $3.8\pm 5.7$ $\textbf{29.5}\pm\textbf{8.2}$ $3.8\pm 5.7$
@@ 5.2 Training Details
6 Experimental Results
6.1 Speech Enhancement Results
The left part of Table 1 presents the speech enhancement results on the VoiceBank-DEMAND test set. Our proposed Ex-Diff model shows strong performance across a range of objective metrics, showcasing significant improvements over USS in PESQ, ESTOI, and SI-SDR, achieving similar performance in SGMSE, and surpassing it in the SI-SIR metric. However, it is worth noting that the combined USS+SGMSE model exhibits a marginally better SI-SDR score compared to Ex-Diff, which can be attributed to two key factors. First, USS as a discriminative model, uses an L1 loss function directly calculated on the predicted and ground truth waveforms, inherently favouring higher SI-SDR scores due to the better (compared to the diffusion training loss) correlation between minimizing waveform distance and maximizing signal-to-interference ratio [25, 2]. Second, the subsequent refinement using the SGMSE diffusion model further contributes to this bias by directly targeting waveform similarity, a factor heavily reflected in SI-SDR. Conversely, Ex-Diff, as an end-to-end generative model, focuses on learning the underlying speech distribution to produce enhanced speech with high perceptual quality. While this approach may not prioritize SI-SDR optimization to the same extent as a concatenated discriminative-diffusion approach, it avoids potential drawbacks of model cascading, such as error propagation and increased complexity. Furthermore, it is crucial to recognize that SI-SDR alone cannot fully encapsulate the nuanced aspects of speech quality, such as naturalness and listener preference, areas where generative approaches like Ex-Diff often excel.
6.2 Vocal Enhancement Results
The MUSDB18 dataset is used for vocal enhancement. According to [28], training models with an appropriate SNR and using incoherent mixing for data augmentation can lead to better performance.
We reprocessed the SNR of the MUSDB18 dataset to 3dB, and adopted an incoherent mixing strategy for the training set, which enhances both the quantity and quality of the data.
The experimental results on MUSDB-mix3dB test set, as shown in Table 1, indicate that our model outperforms all baseline models. It should be noted that, to ensure a fair comparison, our baseline models were also trained from scratch on the MUSDB-mix3dB dataset. In which the USS model achieves good results on the original MUSDB dataset, but performs poorly on the processed MUSDB dataset, which also reflects that the generative model has stronger generalization ability.  [^1] The demo can be accessed at: https://liuzhan22.github.io/ex-diff-demo/..
6.3 Discussions on Selected Examples
@@ 6.4 Ablation Study
Table 2: Results of the ablation study. A single cross-attention layer (as in Fig. 1) performs the best. Removing the attention mechanism by concatenating the latent representation with $\mathbf{x}$ and $\mathbf{y}$, using a full Transformer-like block, or adding up to three attention blocks did not yield better results.
@@ 7 Conclusion
Acknowledgment
This project is funded by the NSFC Young Scientists Fund (Category C) under Grant No. 62501512.